# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 9 Preparatory Results
We continue Example 3.22, and its sequels, Examples 5.55, and 7.74. There exist one path of positive length and two cycles in $AUX[E;F]$.
The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that the following statement holds: for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
@@ 11 Main Cases of the Decision Algorithm
We now define the four main cases of our decision algorithm. We show that we can decide which Main Cases hold in $O(n\,\mathrm{log}\,n)$ time.
@@
    $\Sigma$ is a unary signature. Example 3.17 and its sequels, Examples 5.50, 7.71, 10.95. and 12.109, and 12.114, also Example 3.18 and its sequels, 5.51, 7.72, 10.96, 12.110, and 12.115 illustrate this case.
@@
    Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Example 3.19 and its sequels Examples 5.52, 13.117, further Example 3.20 and its sequels, Examples 5.53, 13.118, exemplify this case.
@@
    $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total. Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132, moreover Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case.
@@
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
@@
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
@@ 12 Main Case 1
In this main case, $\Sigma$ is a unary signature, and $E$ and $F$ are GTESs over $\Sigma$.
@@
Proof.
Lemma 9.83, Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
We present a faster and simpler algorithm for solving our decidability problem. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. Example 3.17 and its sequel, Example 5.50, furthermore Example 3.18 and its sequel, 5.51 illustrate this main case. We assume that we have run Algorithm CAD on $\Sigma$, $E$, and $F$, and study its output, see Proposition 4.47.
@@
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Let $a\in C\langle E\cup F,\emptyset\rangle$ be such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle F$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. Assume that Conditions (a) and (b) hold:
@@
    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
[…]
@@
[…]
@@
Base Case: $n=1$. Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $\delta_{1}[a_{11},\ldots,{k_{1}}]=\sigma(a_{11},\ldots,{k_{1}})$, and $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. Hence by Condition (b), and the rewrite rule above there exists a rewrite rule
$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$
for some $b\in C\langle E,F\rangle$ with $b\subseteq a$. Then we have $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$.
@@
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
@@
  • ($\mathrm{b}^{\prime}$)
[…]
Then we obtain $\delta_{n+1}$ from $\delta_{n}$ replacing the $\ell$th occurrence from left-to right of $\diamondsuit$ with the term $\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$. Consequently,
@@
$\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)}),a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},a_{n,\ell},a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
$a_{n1}=a_{(n+1)1},\ldots,a_{n(\ell-1)}=a_{(n+1)(\ell-1)},a_{n\ell}=a_{n\ell}a_{n(\ell+1)}=a_{(n+1)(\ell-m+1)},\ldots,a_{nk_{n}}=a_{(n+1)k_{(n+1)}}$
@@
$\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)})\dot{\rightarrow}a_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (1)
We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $a$ from a proper extension of $a_{n\ell}$. Therefore by the rewrite rule (1) and Condition ($\mathrm{b}^{\prime}$), there exists a rewrite rule
@@
for some $b_{n\ell}\in C\langle E\cup F,E\rangle$ with $b_{n\ell}\subseteq a_{n\ell}$. Let
$b_{n1}=b_{(n+1)1},\ldots,b_{n(l-1)}=b_{(n+1)(l-1)},\ldots,b_{n(l+1)}=b_{(n+1)(l-m+1)},\ldots,b_{nk_{n}}=b_{(n+1)k_{(n+1)}}$
By the induction hypothesis, there exists $b\in C\langle E,F\rangle$ with $b\subseteq a$ such that the following sequence of reductions holds:
@@
$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{(}n+1)}]=\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},\sigma(b_{(n+1)\ell},\ldots,b_{(n+1)(\ell+m)}),b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},b_{n,\ell},b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]=\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$, ∎
@@
()($\Rightarrow$)( ⇒ ) Let $t\in W$. Then by our assumption, $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$. Assume that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$ is similar. Let $b\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reach $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant $b$. Then there exists $\delta\in C_{\Sigma,1}$ such that $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Consider any rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b$ in $R\langle E\cup F,\emptyset\rangle$. with right-hand side $b$. We now distinguish two cases depending on the value of $m$.
Case 1: $m=1$. Then we consider the rewrite rule $\sigma(b_{1})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$. Then let $c_{1}\in C\langle E,F\rangle$ with $c_{1}\subseteq b_{1}$, and let $t_{1}\in T_{\Sigma}$ such that $t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$. Then we have
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(b_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.

### v2

@@ 9 Preparatory Results
We continue Example 3.23, and its sequels, Examples 5.56, and 7.75. There exist one path of positive length and two cycles in $AUX[E;F]$.
The first cycle: Its length is $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The path of positive length: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The second cycle: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that the following statement holds: for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
  • 1.
    ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
  • 2.
    if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
@@ 11 Main Cases of the Decision Algorithm
We now recall the four main cases of our decision algorithm. We show that we can decide in $O(n\,\mathrm{log}\,n)$ time which main cases hold.
@@
    $\Sigma$ is a unary signature.
@@
    Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
@@
    $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total.
@@
    $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total.
Recall that Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
@@
  • 8.
    $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data;
  • 9.
@@
  • 11.
    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
@@ 12 Main Case 1
In this main case, $\Sigma$ is a unary signature, and $E$ and $F$ are GTESs over $\Sigma$. We define the Algorithm Nested Partial Depth First Search (NPDFS for short) on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$. NPDFS runs in $O(n^{2})$ time. Furthermore, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle)$ if and only if NPDFS outputs true. Thus, we can decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. First we note that we can also apply the decision algorithm of Champavère et al [10].
Lemma 9.93, Proposition 9.94, and Proposition 9.92 imply the following theorem.
@@
We present a faster and simpler algorithm for solving our decidability problem. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48. Example 3.18 and its sequel, Example 5.51, furthermore Example 3.19 and its sequel, 5.52 illustrate this main case.
@@
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Let ${b}\in C\langle E\cup F,\emptyset\rangle$ be such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$. Assume that Conditions (a) and (b) hold:
@@
    where $n\geq 1$ and for every $i\in\{\,1,\ldots,n\,\}$, $\left(k_{i}\in{\mathbb{N}},\;\delta_{i}\in CON_{\Sigma,k_{i}},\mbox{ and }{b}_{i1},\ldots,{b}_{ik_{i}}\in C\langle{E\cup F},\emptyset\rangle\right)$.
@@
    ${c}_{n1},\ldots,{c}_{nk_{n}}\in C\langle E,F\rangle$ such that ${c}_{ni}\subseteq{b}_{ni}$ for every $i\in\{\,1,\ldots,k_{n}\,\}$.
Then there exists ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$ and for every $j\in\{\,1,\ldots,n\,\}$, there exists ${c}_{j1},\ldots,{c}_{jk_{j}}\in C\langle E,F\rangle$ such that ${c}_{ji}\subseteq{b}_{ji}$ for every $i\in\{\,1,\ldots,k_{j}\,\}$, and the following sequence of reductions holds:
$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$,
@@
Base Case: $n=1$. Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $\delta_{1}[{b}_{11},\ldots,{k_{1}}]=\sigma({b}_{11},\ldots,{k_{1}})$, and $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$. Hence by Condition (b), and the rewrite rule above there exists a rewrite rule
$\sigma({c}_{11},\ldots,{c}_{1k_{1}})\dot{\rightarrow}{c}\in R\langle E,F\rangle$
for some ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$. Then we have $\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{c}$.
@@
    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}$,
    where $n\geq 1$ and for every $i\in\{\,1,\ldots,n+1\,\}$, $\left(k_{i}\in{\mathbb{N}},\;\delta_{i}\in CON_{\Sigma,k_{i}},\mbox{ and }{b}_{i1},\ldots,{b}_{ik_{i}}\in C\langle{E\cup F},\emptyset\rangle\right)$.
  • ($\mathrm{b}^{\prime}$)
    ${c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}\in C\langle E,F\rangle$ such that ${c}_{(n+1)i}\subseteq{b}_{(n+1)i}$ for every $i\in\{\,1,\ldots,k_{n+1}\,\}$.
@@
$\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{(n+1)}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(\ell-1)},\sigma({b}_{(n+1)\ell},\ldots,{b}_{(n+1)(\ell+m)}),{b}_{(n+1)(\ell+1)},\ldots{b}_{(n+1)k_{(n+1)}}]$,
$\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(\ell-1)},{b}_{n,\ell},{b}_{(n+1)(\ell+1)},\ldots{b}_{(n+1)k_{(n+1)}}]$,
${b}_{n1}={b}_{(n+1)1},\ldots,{b}_{n(\ell-1)}={b}_{(n+1)(\ell-1)},{b}_{n(\ell+1)}={b}_{(n+1)(\ell-m+1)},\ldots,{b}_{nk_{n}}={b}_{(n+1)k_{(n+1)}}$
@@
We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$. By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reaches b{b} from a proper extension of ${b}_{n\ell}$. Therefore by the rewrite rule (1) and Condition ($\mathrm{b}^{\prime}$), there exists a rewrite rule
$\sigma({c}_{(n+1)\ell},\ldots,{c}_{(n+1)(\ell+m)})\dot{\rightarrow}{c}_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (2)
for some ${c}_{n\ell}\in C\langle E\cup F,E\rangle$ with ${c}_{n\ell}\subseteq{b}_{n\ell}$. Let
${c}_{n1}={c}_{(n+1)1},\ldots,{c}_{n(l-1)}={c}_{(n+1)(l-1)},\ldots,{c}_{n(l+1)}={c}_{(n+1)(l-m+1)},\ldots,{c}_{nk_{n}}={c}_{(n+1)k_{(n+1)}}$
By the induction hypothesis, there exists ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$ such that the following sequence of reductions holds:
$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$,
@@
$\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{(}n+1)}]=\newline \delta_{n}[{c}_{(n+1)1},\ldots,{c}_{(n+1)(\ell-1)},\sigma({c}_{(n+1)\ell},\ldots,{c}_{(n+1)(\ell+m)}),{c}_{(n+1)(\ell+1)},\ldots{c}_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n}[{c}_{(n+1)1},\ldots,{c}_{(n+1)(\ell-1)},{c}_{n,\ell},{c}_{(n+1)(\ell+1)},\ldots{c}_{(n+1)k_{(n+1)}}]=\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$. ∎
@@
$(\Rightarrow)$ Let $t\in W$. Then by our assumption, $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$. Assume that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$ is similar. Let ${c}\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reaches $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant c{c}. Then there exists $\delta\in CON_{\Sigma,1}$ such that $\delta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Consider any rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E\cup F,\emptyset\rangle$. with right-hand side c{c}. We now distinguish two cases depending on the value of $m$.
Case 1: $m=1$. Then we consider the rewrite rule $\sigma({c}_{1})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$. Then let ${d}_{1}\in C\langle E,F\rangle$ with ${d}_{1}\subseteq{c}_{1}$, and let $t_{1}\in T_{\Sigma}$ such that $t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{d}_{1}$. Then we have
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma({c}_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{c}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.